BRICS Notes Series, 1998

July 7, 2003

This document is also available as PostScript and DVI.

References

[NS-98-8] Olivier Danvy and Peter Dybjer, editors. Proceedings of the 1998 APPSEM Workshop on Normalization by Evaluation, NBE ’98 Proceedings, (Gothenburg, Sweden, May 8–9, 1998), December 1998.

[NS-98-7] Abstract, PostScript, PDF, DVI.

John Power. 2-Categories. August 1998. 18 pp.

[NS-98-6] Abstract, PostScript, PDF, DVI.

Carsten Butz, Ulrich Kohlenbach, Sren Riis, and Glynn Winskel, editors. Abstracts of the Workshop on Proof Theory and Complexity, PTAC ’98, (Aarhus, Denmark, August 3–7, 1998), July 1998. vi+16 pp.

[NS-98-5] Abstract, PostScript, PDF.

Hans Httel and Uwe Nestmann, editors. Proceedings of the Workshop on Semantics of Objects as Processes, SOAP ’98, (Aalborg, Denmark, July 18, 1998), June 1998. 50 pp.

[NS-98-4] Abstract, PostScript, PDF.

Tiziana Margaria and Bernhard Steffen, editors. Proceedings of the International Workshop on Software Tools for Technology Transfer, STTT ’98, (Aalborg, Denmark, July 12–13, 1998), June 1998. 86 pp.

[NS-98-3] Abstract, PostScript, PDF.

Nils Klarlund and Anders Mller. MONA Version 1.2 — User Manual. June 1998. 60 pp.

[NS-98-2] Peter D. Mosses and Uffe H. Engberg, editors. Proceedings of the Workshop on Applicability of Formal Methods, AFM ’98, (Aarhus, Denmark, June 2, 1998), June 1998. 94 pp.

[NS-98-1] Olivier Danvy and Peter Dybjer, editors. Preliminary Proceedings of the 1998 APPSEM Workshop on Normalization by Evaluation, NBE ’98, (Gothenburg, Sweden, May 8–9, 1998), May 1998.